OpenStudy (unknownunknown):

Is this another error, or is there something I'm missing? For question 3, why not just use partial dw / partial x = (ze^y + e^z) ? Why is it setting values to 0 such as in (0 + e0) ? What has that answer got to do with anything? Thanks.

OpenStudy (michele_laino):

since your professor, want to evaluate \[\Large \frac{{\partial w}}{{\partial x}}\] along a variety, which lies on the plane xy, and the cartesian equation of the xy-plane is \[\Large z = 0\]

OpenStudy (australopithecus):

I am pretty sure it is because there is no z term in the constraint, thus z = 0 also by taking the partial derivative in terms of z of the constraint you get dz=0

OpenStudy (michele_laino):

the equation of that variety is: \[\Large {x^2}y + x{y^2} = 1\]

OpenStudy (australopithecus):

From question one you know what dw equals so you can apply the constraint and write it in terms of the partial derivative in respect to x

OpenStudy (unknownunknown):

And then how does the third part equal the fourth part? From dx + 2xy+y^2/(x^2+2xy) = (x^2+4xy+y^2)/(x^2+2xy)

myininaya (myininaya):

http://www.mit.edu/~hlb/tempdir/done/MIT18_02SC_pb_42_comb.pdf take a look at this

myininaya (myininaya):

this seems like basically the same pdf but they have z=0

myininaya (myininaya):

If that link didn't work.

OpenStudy (unknownunknown):

Wow @myininaya, that's great, how did you find that? I'd say 50% of the pdfs in this course have obvious things like that omitted

OpenStudy (australopithecus):

There is no z term in the constraint, therefore z = 0, the partial derivative of the constraint in respect to z is equal to 0 so you sub that into the equation you wrote for dw in the first question

OpenStudy (australopithecus):

(xe^y + xe^z + ye^z)(0) = 0 because dz = 0

OpenStudy (australopithecus):

but yeah you can see that on that pdf that was posted

OpenStudy (unknownunknown):

And then how does the third part equal the fourth part, I can see the numerator being x^2+4xy+y^2, but the denominator there wouldn't make sense solving for dx?

OpenStudy (australopithecus):

you factor out the dx

OpenStudy (australopithecus):

then you just divide both sides by dx and you get dw/dx

myininaya (myininaya):

\[dw=1 dx+\frac{2xy+y^2}{x^2+2xy} dx \\ dw=\frac{x^2+2xy}{x^2+2xy} dx+\frac{2xy+y^2}{x^2+2xy} dx \\ dw=(\frac{x^2+2xy}{x^2+2xy}+\frac{2xy+y^2}{x^2+2xy} )dx\] what he said you can combine the fractions now

OpenStudy (australopithecus):

When taking the total derivative of the constraint you can find what dy equals then you can sub that into the equation from the first question so you have it only in terms of dx and dw

OpenStudy (unknownunknown):

Ahh yep I see it. Makes sense now. But that z=0 not being included, and the third term not being cancelled in the original pdf, those errors are rampant in this course. Yet the pdf you posted myinanaya seems without errors. How did you locate that? I wonder if they exist for all the other pdfs..

OpenStudy (australopithecus):

That isn't really an error, you can infer it from the function

myininaya (myininaya):

I just googled part of the question.

OpenStudy (unknownunknown):

Seems so. But at this level this topic is entirely new to me, so the explanation there that z=0 would have been crucial.

OpenStudy (australopithecus):

it can help to look at functions on wolfram alpha

OpenStudy (australopithecus):

most calc text books have a list of multivariable formulas and their specific shapes

myininaya (myininaya):

this is the exact thing I typed into google "3. Now suppose w is as above and x2y + y2x = 1. Assuming x is the independent variable," and it was the 3rd pdf/link thingy

OpenStudy (australopithecus):

yeah most questions asked by profs are online if you dig deep enough or you can find ones similar

myininaya (myininaya):

There are some people who are way more expert than me at being an expert googler

OpenStudy (unknownunknown):

I can see now the z=0 term isn't necessary, but the thing is is that most of the pdf's actually contain proper errors. I've posted quite a few of them in the Multivar MIT section here. It seems strange there is a correct pdf here, why they wouldn't include it in the actual course site.

OpenStudy (michele_laino):

more precisely, the variety which is considered by your teacher is defined by the subsequent cartesian equations: \[\Large \left\{ \begin{gathered} {x^2}y + x{y^2} = 1 \hfill \\ z = 0 \hfill \\ \end{gathered} \right.\] @unknownunknown

OpenStudy (unknownunknown):

Actually not sure I understand entirely the inference if z=0 isn't specified. If it isn't specified in the constrain, this could mean z would be a plane and could contain all values of z no?

OpenStudy (unknownunknown):

Eg: constraint x=3, this would still allow for all values of x and y.

OpenStudy (unknownunknown):

z and y*

OpenStudy (australopithecus):

z = 0, because it is not present in the equation. This constraint is restricted to the x, y plane

OpenStudy (unknownunknown):

A function in the xy plane right? Which would extend to + and - infinity in the z plane with that shape.

OpenStudy (unknownunknown):

Taking that constraint, if I make z = 5, or z = 119, or anything, it is still a valid constraint. My z value can be anything.

OpenStudy (unknownunknown):

The constraint will still hold.

OpenStudy (australopithecus):

If there was a z variable present in it but there isn't, its an equation of 2 variables

OpenStudy (unknownunknown):

But our main function does have a z variable, so as long as it intersects in the x & y point, the z can be anything no?

OpenStudy (australopithecus):

no the constraint restricts it to the x, y plane

OpenStudy (australopithecus):

by the constraint z=0 and dw/dz = 0

OpenStudy (unknownunknown):

If it's specified as z=0 yes, but in the original pdf it isn't right? Are you sure we can infer that if it isn't specified? For example imagine any function in 3 space, and we restrict it to x=3. At x=3 we should still have a plane of all values in the zy plane located at x=3 no?

OpenStudy (australopithecus):

The original pdf is right, because the constraint is confined to z=0, or specifically the xy plane just like the equation x^2 + y^2 = 1 is constrained to the x, y plane

OpenStudy (australopithecus):

It can be inferred from the constraint that z = 0

OpenStudy (australopithecus):

Although, I do agree it might have been nice if they showed it more so in the answer.

OpenStudy (unknownunknown):

So where in the original pdf is the constraint z=0?

OpenStudy (australopithecus):

Look at the graph of the constraint http://www.wolframalpha.com/input/?i=xy^2+%2B+xy+%3D+1+graph

OpenStudy (australopithecus):

It is confined to the x,y plane

OpenStudy (australopithecus):

since the constraint is confined to the x,y plane z =0

OpenStudy (unknownunknown):

But we can set z to be anything in that graph, and it would extend into space, and still be correct.

OpenStudy (australopithecus):

No the constraint forbids that, because you are looking at the slope in terms of the constraint and the constraint doesn't ever leave the x,y plane thus z always equals 0

OpenStudy (australopithecus):

the constraint restricts what you are looking at

OpenStudy (australopithecus):

The constraint says, dw/dz = 0 it also says z = 0

OpenStudy (australopithecus):

to ignore that is to ignore the constraint

OpenStudy (australopithecus):

The question is asking for a slope in terms of the constraint so to ignore the constraint would give you an incorrect answer

OpenStudy (unknownunknown):

Hmm, so we're not just constraining the x and y variables in w? We must also then assume z=0?

OpenStudy (australopithecus):

We are constraining all variables, we are not assuming z = 0, in terms of this constraint z always equals 0 because it is a two variable equation. There are no assumptions.

OpenStudy (australopithecus):

Sorry wrong photo, here this can be helpful for identifying multivariate functions

OpenStudy (unknownunknown):

So how is the constraint different say, from setting z=3 as a level and viewing the function as a contout plot? This would then give the xy plane at z=3. A constraint of x=3 wouldn't be a level of the zy plane? Are they two different concepts?

myininaya (myininaya):

hey @Australopithecus I haven't been saying anything because I have been thinking about this constraint problem for awhile now. Let me see if you agree. The constraint x^2y+y^2x=1 shows that z doesn't depend on x and only y does. So z can be anything that isn't related to x. And if z is not related to x then it can't be related to y since y and x share a relationship. so dz/dx=0 but I don't have a good reason yet to conclude z=0 dz/dx=0 implies z=constant but I don't know how to explain why the constant would be zero sorry not trying to disagree or anything I'm still trying to process why they didn't specify that z=0 in the pdf

OpenStudy (michele_laino):

z=0, is a work hypothesis @myininaya

OpenStudy (australopithecus):

Yes you are right sorry, I think they just picked z = 0 arbitrarily, sorry you are right it goes +/- z axis infinitely

OpenStudy (australopithecus):

I apologize I made a mistake, good this is cleared up now though

OpenStudy (unknownunknown):

Ahh interesting, thanks though @Australopithecus for the insights. So can we conclude the original pdf contains an error by omitting that z=0?

OpenStudy (australopithecus):

No we cannot conclude that

OpenStudy (australopithecus):

They just assumed you would pick z = 0, but you could pick any z value

OpenStudy (unknownunknown):

But, the correct form would be as it is in the first line.. dw = (ze^y + e^z)dx, since this term allows us to pick anything. But then in the next line, it picks 0 without specifying it as such. The first line would allow us to pick anything, the second assumes 0, and it gives an equality there. That is not an error?

OpenStudy (australopithecus):

they didn't need to specify because in terms of the constraint, dz = 0

OpenStudy (australopithecus):

so you could sub z into it but it would all become zero anyways so why would you

OpenStudy (unknownunknown):

So it's asking to find \[\partial dw / \partial dx\] , by definition the coefficient of dx must be that partial. Not (0 + 1) as it says.

OpenStudy (australopithecus):

no no, I am just using d instead of daron because I am lazy dont get confused

OpenStudy (australopithecus):

∂z = 0

OpenStudy (australopithecus):

the partial derivative of the constraint in terms of z is ∂w/∂z = 0 therefore ∂z =0 thus, (xe^y + xe^z + ye^z)(0) = 0 because ∂z = 0 so you could sub a z value into that but it would still become equal to 0

OpenStudy (unknownunknown):

Yeah that's fine for the third term. But looking at the first term where we have dx. In the first line that is correct, as it is partial w/ partial x, yet in the second line it spontaneously changes to 1?

OpenStudy (australopithecus):

e^0 = 1

OpenStudy (unknownunknown):

Yes, but why would you even put e^0?

OpenStudy (australopithecus):

because the answerer chose z = 0

OpenStudy (unknownunknown):

But we're then evaluating a partial, when its not asking that.

OpenStudy (australopithecus):

like you were saying you could look at this slope at different values of z

OpenStudy (unknownunknown):

Say we take the derivative of some f(x) = x^2, we get 2x. We can't then say f'(x) = 0 because we decided to arbitrarily evaluate the derivative at 0.

OpenStudy (australopithecus):

z is a constant, in your example x is not a constant

OpenStudy (australopithecus):

the constraint is a two variable equation

OpenStudy (australopithecus):

thus z is not a variable it is a constant

OpenStudy (unknownunknown):

z isn't a constant, still a variable in dw / dx. We can input any value of z right.

OpenStudy (australopithecus):

no because dw/dz = 0 according to the constraint

OpenStudy (australopithecus):

that is the derivative of a constant

OpenStudy (australopithecus):

thus z is a constant

OpenStudy (australopithecus):

so you have to pick one

OpenStudy (unknownunknown):

Okay so going back to that then, assuming dw/dz = 0, and z is a constant. Why do we make that assumption again?

OpenStudy (australopithecus):

It is not an assumption, take the partial derivative of: \[x^2y + y^2x = 1\] in terms of ∂w/∂z what do you get?

OpenStudy (australopithecus):

you get ∂w/∂z = 0 hence the constraint says that z is a constant, its slope does not change, therefore it is not a variabl

OpenStudy (unknownunknown):

So then your first graph would be correct instead of the second graph? Since in the second z isn't a constant but taking many different values.

OpenStudy (australopithecus):

They are both actually the same graph, just one is in 2D. You could look at that graph at any level of z and it would look the same though. If z was a variable in the constraint this would not be the case. It is the same deal with this derivative

OpenStudy (unknownunknown):

If you wanted to allow all values of z in the constraint, how would you then express that?

OpenStudy (australopithecus):

you are looking at a larger three variable function constrained by a two variable function

OpenStudy (australopithecus):

The derivative will just change by a constant value